If A is the number of people in the sample who preferred Product A, using your data, what is xA bar and sA (Product A: 13, Product B: 4, Product C: 13)
Answer #1

Solution:

n=A+B+C=13+4+13=30

X bar=sample proportion=x/n=13/30=0.4333333

sA =sample standard deviation=sqrt(p^*(1-p^)/n

=sqrt(0.4333333*(1-0.4333333)/30)

= 0.09047201

ANSWERS IN BOLD:

0.4333333

0.09047201